An object, with mass 93 kg and speed 23 m/s relative to an observer, explodes into two pieces, one 3 times as massive as the other; the explosion takes place in deep space. The less massive piece stops relative to the observer. How much kinetic energy is added to the system during the explosion, as measured in the observer's reference frame
Answer #1

Here,

v= 23 m/s

Then,

Solution; mi - Here, The mass is, 3 x 93 kg 4 = 179 279 kg x 93kg 93 kg v = 23 mis M2 = 4 - ma 4 )) V2 = 0 Now, using conservThen, The initial kinetic energy is, - k. Ei = 1 (mit m2) v2 2 - K.Ei - 0.5 05/292 +)*(23) = KFi = 0.51 69.75 +23-25) X(23)?